OpenStudy (anonymous):

A professor asked the students in his class to rate the guest lecturer. The lecturer received a rating of Excellent from 7 of the students, while 14 students rated the lecturer as Average, and 4 believed the lecturer was Poor. What percent of the students thought the lecturer was Excellent?

OpenStudy (ganpat):

percentage would be = (No of students rated excellent) / (Total number of students) * 100 So, Percent = 7 / (7 + 14 + 4) * 100 = 7/ 25 * 100 = 28 %...

OpenStudy (ganpat):

is that a wrong answer ??

OpenStudy (anonymous):

no, thank you so much !

OpenStudy (ganpat):

wc..